Naruko885:

Ada says that the length of diagonal SQ is two times the length of diagonal OM. Is Ada correct? Justify your answer and show all your work. Your work should state the theorem you used to find the lengths of the diagonals.

Angle:

I cannot access the picture

Angle:

Assuming that LMNO is a square we can try to work this problem out by using the Pythagorean theorem in order to find the length of SQ and the length of OM

Angle:



Angle:

the Pythagorean theorem is \(a^2 + b^2 = c^2 \) where a and b are sides of the rectangles (or "legs") and c is the "hypotenuse" or SQ because of this, we can say that \(14^2 + 7^2 = (SQ)^2\) and try to solve for or estimate SQ

Angle:

meanwhile

Angle:

using the same formula I described above... \(7^2 + 7^2 = (OM)^2\) This should be enough information to get you to your answer :) If you still need help, please do reply @Naruko885 ^_^ I would really appreciate hearing where you're having trouble so I can help you better :3

Naruko885:

I think I get what your saying but can you help me like plug in the numbers in the right spots? thanks.

Angle:

I already plugged in all the numbers for you ;) \(14^2 + 7^2 = (SQ)^2\) and \(7^2 + 7^2 = (OM)^2 \) all you have to do is solve for SQ and OM

Naruko885:

SQ 98 and OM 49?

Naruko885:

oh wait I think I did it wrong.

Naruko885:

wait did I do it wrong, im confused??

Angle:

mhmm something might be wrong \(14^2\) = 196 \(7^2\) = 49 196 + 49 = 245

Angle:

then the square root of 245 is about 15.6

Naruko885:

ohhh I accidently x not + sorry. I get it now.

Naruko885:

thank you so much!